Please solve the following inequalities and write the solutions in interval notation.
x^3-2x^2

Answer by dolly(163)


x^3-2x^2 ==> x^2(x-2)< or = 25(x-2)
==> x^2 < or = 25 [dividing by (x-2) on both the sides]
==> x < or = 5, - 5
The solution is [- 5, 5]